Grade 11Mechanics

hi my name is jenil i am in std 11 i have one dout numerical no one is solving it. a man going with velocity 5 m/s.no other force acting so the work done is?

jenil
6 Years agoGrade 11

4 Answers

Kaushal Negi
6 Years ago
Okay Jenil read carefully
Since the acceleration of the man is zero because he is moving with a constant velocity, i.e. 5m/s.
Work done = F.s
F=m*a
As acceleration is zero so the force will also be zero.
Hence work done by the man will be zero
Vikas TU
6 Years ago
Dear student 
You have mentioned in the question  that the velocity is constant , if it wil vary with some another variabe , Then it would be in terms of some variables,
acceleration = dv /dt 
So a = 0 
So work done wil be zero.
Good Luck 
aswanth nayak
6 Years ago
Dear Student,
 
Youhave mentioned in the question that the velocity is constant, if it will vary with some other variable,then it would be in terms of some variables, Since the acceleration of man is zero because he is moving with a constant velocity
Therefore , W = F.S
and we know that F=m*a
As Acceleration is zero, Force will also be zero
Hence here work done by the man will be zero.
 
Hope this helps you
Regards
 
 
Ayush_Deep
6 Years ago
Dear friend , since there is no force acting on body ,there will be no work done
Work done = F*S
F=0 so W=0
Thanks